OpenStudy (anonymous):

How much mass is equivalent to 1 MJ of energy?

OpenStudy (anonymous):

A. 1.1 10^-5 kg B. 1.1 10^-11 kg C. 3.3 10^-3 kg D. 9.0 10^22 kg

OpenStudy (cwrw238):

e = mc^2

OpenStudy (anonymous):

can you help me to use it?

OpenStudy (anonymous):

cause i got this question and have never been taught to put it into equation

OpenStudy (cwrw238):

c = speed of light in meters / second

OpenStudy (cwrw238):

Einstein's famous equation

OpenStudy (anonymous):

still confused what do i do with that to tell how many kg its equal to?

OpenStudy (cwrw238):

speed of light is 3 * 10^8 m / s

OpenStudy (cwrw238):

so m = 1,000,000 / (3*10^8)^2 = 1.1 8 10^-11 kg

OpenStudy (cwrw238):

* 1.1 * 10^-11 kg

OpenStudy (cwrw238):

e = m c^2 so m = e / c^2

OpenStudy (cwrw238):

mass = energy / square of speed of light

OpenStudy (cwrw238):

an amazing amount of energy is formed from such a tiny mass. That's why the atomic and hydrogen bombs are so destructive.

OpenStudy (anonymous):

where di the 1,000,000 come from?

OpenStudy (cwrw238):

1 MJ = 1 mega joule or 1,000,000 joules of energy

OpenStudy (cwrw238):

written in scientific notation it is 1 * 10^6